change in the demand of cold drinks from 700 units to 1000 units because of the fall in the price from Rs 15 to Rs 10. Calculate the price elasticity of cold drinks


1
Expert's answer
2021-05-11T15:13:32-0400

price elasticity"=\frac {Q_2-Q_1 }{(Q_2+Q_1)/2}\div\frac{P_2-P_1}{(P_2+P_1)/2}"


Change in quantity "= \dfrac{1000-700}{\dfrac{700+1000}{2}}"

"= \frac{300}{850}=0.3529"

Change in price "= \dfrac{10-15}{\dfrac{10+15}{2}}" 

"= \dfrac{-5}{12.5} =0.4" 


Hence, Price elasticity of demand "= \dfrac{0.3529}{-0.4}=-0.88"


"= 0.88"